OpenStudy (anonymous):

4n – 28 = 8 0 5 7 none of the above

OpenStudy (anonymous):

Its 5

OpenStudy (anonymous):

all you have to do is this \[4\left( 5 \right)-28=8\]

OpenStudy (anonymous):

Explain to this person HOW to do it :) @Lover9

OpenStudy (anonymous):

i did :)

OpenStudy (anonymous):

\[4\left( 5 \right)=20\]

OpenStudy (anonymous):

wait no.....im doing it wrong hahaha

OpenStudy (anonymous):

its D sorry :)

OpenStudy (anonymous):

\[4\left( 7 \right)=28 \] \[28-28=0\]

OpenStudy (anonymous):

Could i maybe get a medal plz ? :)

OpenStudy (anonymous):

c makes it go to 0 and a is pretty obvious thats wrong so its D.none of the above